# Monotone Sequences of Metric Spaces with Compact Limits

### v1

@@ 4.5. Integral Current Spaces
In [SW11], Sormani and Wenger define an $m$-dimensional integral current space, $(X,d,T)$, as a metric space, $(X,d)$, with an integral current, $T$, defined on the closure of $(X,d_{X})$ such that the $\operatorname{set}(T)=X$ where
@@
Suppose the sphere ${\mathbb{S}}^{2}$ is endowed with the standard round metric, $g_{{\mathbb{S}}^{2}}$, and the singular metric, $g_{h}$,
@@
defined on a sphere parametrized by $r\in[0,\pi]$ and $\theta\in[0,2\pi]$, where $h(r(x))$ is smooth on $r^{-1}(0,\pi)$, increasing on $[0,\pi/2]$ and decreasing on $[\pi/2,\pi]$ with $h(r)=h(\pi-r)$, but only continuous at the poles $p_{0},p_{\pi}$ where $r=0,\pi$ respectively. We can define a unique compact metric space, $({\mathbb{S}}^{2},d_{h})$, using $g_{h}$-lengths of curves, $C$, such that
@@
and where $T_{h}=[[{\mathbb{S}}^{2}]]$ in the sense of Definition 4.9. So $\operatorname{spt}_{d_{h}}(T_{h})={\mathbb{S}}^{2}$ and $\partial T_{h}=0$. The poles are included in $\operatorname{set}_{d_{h}}(T_{h})$ iff
@@
We see that $T_{h}=[[{\mathbb{S}}^{2}]]$ in the sense of Definition 4.9 because
@@
then $({\mathbb{S}}^{2},d_{cusp})$ is a compact metric space with cusp singularities at the poles, and the triple $(\operatorname{set}_{d_{cusp}}([[{\mathbb{S}}^{2}]],d_{h},[[{\mathbb{S}}^{2}]])$ is an integral current space with
@@
(4.62) $|\nabla(h_{j}\circ r)|=|h_{j}^{\prime}(r)|\leq|h_{j+1}^{\prime}(r)|\leq|h_{cusp}^{\prime}(r)|.$
Thus we have a monotone increasing sequence of metric tensors as described in Remark 2.4. We have uniformly bounded radial arclength parameters. The radial arclengths,
@@ 4.6. Intrinsic Flat Distance
In [SW11], Sormani-Wenger prove this implies there is a common complete metric space, $(Z,d_{Z})$, and distance preserving maps, $f_{j}:X_{j}\to Z$ such that, $T_{j}=f_{j\#}S_{j}$ converge weakly to $T_{\infty}=f_{\infty\#}S_{\infty}$ as integral currents on $Z$. Thus by Ambrosio-Kirchheim semicontinuity of mass,
@@ 4.7. Intrinsic Flat Convergence Theorem
where $T_{\infty}=T$ in the sense defined in Definition 4.9 and
@@
so we have volume preserving SWIF convergence.
@@
  • Monotonicity of metric tensors in (1.1) implies monotonicity of distances as in Definition 2.2 by Lemma 2.3 as in Remark 2.4.
@@
Under the hypotheses Theorem 4.16, we have the metric space $(Z,d_{Z})$ as constructed in Proposition 3.9 depicted in Figure 4.3 with distance preserving maps,
@@
Furthermore, there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, $T_{\infty}$, whose support lies on the central spine in $Z$,
(4.83) $\operatorname{spt}(T_{\infty})=Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\})\subset\zeta_{j}(Z_{j,\infty})\subset Z.$
Finally $T=\chi_{\infty\#}^{-1}T_{\infty}$ is an integral current on $(X,d_{\infty})$ as in Definition 4.9, and
(4.84) ${\mathbf{M}}_{d_{\infty}}(T)=\lim_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
Note that before we prove this proposition, we do not yet know that $T$ is an integral current on $(X,d_{\infty})$. See Remark 4.8. After we prove this proposition we will apply it to prove the same statement holds without needing a subsequence in Corollary 4.18.
@@
Since $(Z,d_{Z})$ is compact, we can apply the Ambrosio-Kirchheim Compactness Theorem, Theorem 4.10, to conclude that there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, $T_{\infty}$.
Figure 4.4. Here we see integral currents $T_{j}$ in $Z$ and $B_{d_{Z}}(z,t)$ of the proof that $\operatorname{spt}(T_{\infty})\subset Z_{0}=\chi_{\infty}(X)\subset Z$ as claimed in (4.85).
@@
(4.85) $\operatorname{spt}(T_{\infty})\subset Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\}\subset Z_{j,\infty})\subset Z.$
@@
(4.88) $||T_{\infty}||(B_{d_{Z}}(z,t))\leq\liminf_{j\to\infty}||T_{j}||(B_{d_{Z}}(z,t))=0$
so $z\notin\operatorname{spt}(T_{\infty})$. Thus $\operatorname{spt}(T)\subset Z_{0}$.
@@
(4.89) $S=\chi^{-1}_{\infty\#}T_{\infty}$
@@
We claim that $S=T$ in the sense of Definition 4.9. Given any tuple, $(\pi_{0},\pi_{1},...,\pi_{m})$ of Lipschitz functions on $(X,d_{j})$, we need only show
(4.90) $T(\pi_{0},\pi_{1},...,\pi_{m})=S(\pi_{0},\pi_{1},...,\pi_{m}).$
By Lemma 4.1 these $\pi_{i}$ are also tuples of Lipschitz functions on $(X,d_{\infty})$, so $S(\pi_{0},...,\pi_{m})$ is well defined.
@@
we have by definition of pushforward that
(4.93) $S(\pi_{0},\pi_{1},...,\pi_{m})$ $=$ $(\chi^{-1}_{\infty\#}T_{\infty})(\tilde{\pi}_{0}\circ\chi_{\infty},\tilde{\pi}_{1}\circ\chi_{\infty},...,\tilde{\pi}_{m}\circ\chi_{\infty})$
(4.94) $=$ $T_{\infty}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
@@
So $T_{j_{k}}$ converge weakly as currents to $T_{\infty}$ in $(W_{j},d_{Z})$: for any Lipschitz tuple, $\omega$, on $(W_{j},d_{Z})$, we have
(4.96) $T_{j_{k}}(\omega)\to T_{\infty}(\omega).$
Thus we have
(4.97) $S(\pi_{0},\pi_{1},...,\pi_{m})=\lim_{k\to\infty}T_{j_{k}}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
@@
Substituting this into (4.97), and taking the limit $j_{k}\to\infty$, we reach our claim in (4.90). We conclude that $S=T$ in the sense of Definition 4.9.
Now we claim $\operatorname{spt}_{d_{Z}}(T_{\infty})=Z_{0}$. Since we know $T_{\infty}=\chi_{\infty\#}T$, this is the same as showing $\operatorname{spt}_{d_{\infty}}(T)=X$. Given any $x\in X$ and any $r>0$ we have
(4.101) $||T||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||T||_{d_{j}}(B_{d_{j}}(x,r))$
@@
(4.102) $B_{d_{j}}(x,r)\supset B_{d_{j}}(x,r-\epsilon_{j}).$
Combining this with the above,
(4.103) $||T||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||T||_{d_{j}}(B_{d_{j}}(x,r-\epsilon_{j}))>0.$
Thus $\operatorname{spt}_{d_{\infty}}(T)=X$ and $\operatorname{spt}_{d_{Z}}(T_{\infty})=Z_{0}$.
Finally, by Ambrosio-Kircheim’s semicontinuity of mass combined with the fact that $\zeta_{\infty}$ is distance preserving we have
(4.104) ${\mathbf{M}}_{d_{\infty}}(T)={\mathbf{M}}_{d_{Z}}(T_{\infty})\leq\liminf_{k\to\infty}{\mathbf{M}}_{d_{Z}}(T_{j_{k}})=\liminf_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
Yet by Lemma 4.7, we know
(4.105) ${\mathbf{M}}_{d_{\infty}}(T)\geq{\mathbf{M}}_{d_{j_{k}}}(T)$
so we have mass convergence. ∎
Corollary 4.18.
Under the hypotheses of Proposition 4.17, we have the conclusions $T_{j}=\chi_{j\#}T$ converge weakly to $T_{\infty}=\chi_{\infty\#}T$ as integral currents in $(Z,d_{Z})$ and
(4.106) ${\mathbf{M}}_{d_{\infty}}(T)=\lim_{j\to\infty}{\mathbf{M}}_{d_{j}}(T).$
without taking a subsequence.

### v2

@@ 4.5. Integral Current Spaces
In [SW11], Sormani and Wenger define an $m$-dimensional integral current space, $(X,d,T)$, as a metric space, $(X,d)$, with an integral current, $T$, defined on the closure of $(X,d)$ such that the $\operatorname{set}(T)=X$ where
@@
Suppose the sphere ${\mathbb{S}}^{2}$ is endowed with the standard round metric, $g_{{\mathbb{S}}^{2}}=dr^{2}+\sin^{2}(r)\,d\theta^{2}$, defined on a sphere parametrized by $r\in[0,\pi]$ and $\theta\in[0,2\pi]$. We define the possibly singular metric, $g_{h}$, induced by a graph, $h\circ r:{\mathbb{S}}^{2}\to[0,\infty)$ so that
@@
where $h(r(x))$ is smooth on $r^{-1}(0,\pi)$, increasing on $[0,\pi/2]$ and decreasing on $[\pi/2,\pi]$ with $h(r)=h(\pi-r)$, but possibly only continuous at the poles $p_{0},p_{\pi}$ where $r=0,\pi$ respectively. We can define a unique compact metric space, $({\mathbb{S}}^{2},d_{h})$, using $g_{h}$-lengths of curves, $C$, such that
@@
and where “$T_{h}=[[{\mathbb{S}}^{2}]]$ viewed as an integral current on $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$” in the sense of Definition 4.9.
So $\operatorname{spt}_{d_{h}}(T_{h})={\mathbb{S}}^{2}$ and $\partial T_{h}=0$. The poles are included in $\operatorname{set}_{d_{h}}(T_{h})$ iff
@@
We see that “$T_{h}=[[{\mathbb{S}}^{2}]]$ viewed as an integral current on $({\mathbb{S}}^{2},d_{{\mathbb{S}}^{2}})$” in the sense of Definition 4.9 because
@@
then $({\mathbb{S}}^{2},d_{cusp})$ is a compact metric space with cusp singularities at the poles, and the triple $(\operatorname{set}_{d_{cusp}}([[{\mathbb{S}}^{2}]]),d_{h},[[{\mathbb{S}}^{2}]])$ is an integral current space with
@@
(4.64) $|\nabla(h_{j}\circ r)|=|h_{j}^{\prime}(r)|\leq|h_{j+1}^{\prime}(r)|=|\nabla(h_{j+1})|\leq|h_{cusp}^{\prime}(r)|$
and $\nabla(h_{j}\circ r)$ has the same direction as $\nabla r$ for all $j\in\mathbb{N}$. Thus we have a monotone increasing sequence of metric tensors as described in Lemma 2.4.
The radial arclengths,
@@ 4.6. Intrinsic Flat Distance
In [SW11], Sormani-Wenger prove this implies there is a common complete metric space, $(Z,d_{Z})$, and distance preserving maps, $f_{j}:X_{j}\to Z$ such that, $T_{j}=f_{j\#}S_{j}$ converge weakly to “$T_{\infty}=f_{\infty\#}S_{\infty}$ as integral currents on $(Z,d_{Z})$”. Thus by Ambrosio-Kirchheim semicontinuity of mass,
@@ 4.7. Intrinsic Flat Convergence Theorem
where “$T_{\infty}=T$ viewed as an integral current on $(X,d_{j})$” in the sense defined in Definition 4.9 and
@@
so we have volume preserving intrinsic flat (${\mathcal{VF}}$) convergence.
@@
  • Monotonicity of metric tensors in (1.1) implies monotonicity of distances as in Definition 2.2 by Lemma 2.3.
@@
Under the hypotheses of Theorem 4.16, we have the metric space $(Z,d_{Z})$ as constructed in Proposition 3.9 depicted in Figure 4.3 with distance preserving maps,
@@
Furthermore, there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, $\hat{T}_{\infty}$, whose support lies on the central spine in $Z$,
(4.85) $\operatorname{spt}(\hat{T}_{\infty})=Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\})\subset\zeta_{j}(Z_{j,\infty})\subset Z.$
Taking $T_{\infty}=\chi_{\infty\#}^{-1}\hat{T}_{\infty}$ we have “$T_{\infty}=T$ viewed as an integral current on $(X,d_{j})$ as in Definition 4.9, and
(4.86) ${\mathbf{M}}_{d_{\infty}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})=\lim_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
In the next section we will apply this proposition to prove SWIF convergence holds without needing a subsequence.
@@
Since $(Z,d_{Z})$ is compact by Proposition  3.9, we can apply the Ambrosio-Kirchheim Compactness Theorem, Theorem 4.10, to conclude that there is a subsequence $T_{j_{k}}$ which converges weakly as currents in $(Z,d_{Z})$ to an integral current, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}$.
Figure 4.4. Here we see integral currents $T_{j}$ in $Z$ and $B_{d_{Z}}(z,t)$ of the proof that $\operatorname{spt}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})\subset Z_{0}=\chi_{\infty}(X)\subset Z$ as claimed in (4.87).
@@
(4.87) $\operatorname{spt}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})\subset Z_{0}=\chi_{\infty}(X)=\zeta_{j}(X\times\{0\}\subset Z_{j,\infty})\subset Z.$
@@
(4.90) $||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}||(B_{d_{Z}}(z,t))\leq\liminf_{j\to\infty}||T_{j}||(B_{d_{Z}}(z,t))=0$
so $z\notin\operatorname{spt}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})$. Thus $\operatorname{spt}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})\subset Z_{0}$.
@@
(4.91) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}=\chi^{-1}_{\infty\#}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}$
@@
We claim that for each $j\in{\mathbb{N}}$, “${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}=T$ viewed as an integral current on $(X,d_{j})$” in the sense of Definition 4.9. Given any tuple, $(\pi_{0},\pi_{1},...,\pi_{m})$ of Lipschitz functions on $(X,d_{j})$, we need only show
(4.92) $T(\pi_{0},\pi_{1},...,\pi_{m})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}(\pi_{0},\pi_{1},...,\pi_{m}).$
First note that by Lemma 4.1 these $\pi_{i}$ are also tuples of Lipschitz functions on $(X,d_{\infty})$, so ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}(\pi_{0},...,\pi_{m})$ is well defined.
@@
we have by definition of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}$ and then the definition of pushforward that
(4.95) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}(\pi_{0},\pi_{1},...,\pi_{m})$ $=$ $(\chi^{-1}_{\infty\#}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})(\tilde{\pi}_{0}\circ\chi_{\infty},\tilde{\pi}_{1}\circ\chi_{\infty},...,\tilde{\pi}_{m}\circ\chi_{\infty})$
(4.96) $=$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
@@
So $T_{j_{k}}$ converge weakly as currents to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}$ in $(W_{j},d_{Z})$. This means, for any Lipschitz tuple, $\omega$, on $(W_{j},d_{Z})$, we have
(4.98) $T_{j_{k}}(\omega)\to{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}(\omega).$
Applying this to $\omega=(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m})$, we have
(4.99) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}(\pi_{0},\pi_{1},...,\pi_{m})=\lim_{k\to\infty}T_{j_{k}}(\tilde{\pi}_{0},\tilde{\pi}_{1},...,\tilde{\pi}_{m}).$
@@
Substituting this into (4.99), and taking the limit $j_{k}\to\infty$, we reach our claim in (4.92). We conclude that “${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}=T$ viewed as an integral current on $(X,d_{j})$” in the sense of Definition 4.9.
Now we claim $\operatorname{spt}_{d_{Z}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})=Z_{0}$. Since we know ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}}=\chi_{\infty\#}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}$, this is the same as showing $\operatorname{spt}_{d_{\infty}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})=X$. Given any $x\in X$ and any $r>0$ we have
(4.103) $||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}||_{d_{j}}(B_{d_{\infty}}(x,r))$
@@
(4.104) $B_{d_{\infty}}(x,r)\supset B_{d_{j}}(x,r/2).$
Combining this with the above, and then the fact that “${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}=T$ as integral currents on $(X,d_{j})$”, we have
(4.105) $||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}||_{d_{\infty}}(B_{d_{\infty}}(x,r))\geq||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}||_{d_{j}}(B_{d_{j}}(x,r/2))=||T||_{d_{j}}(B_{d_{j}}(x,r/2))>0.$
Thus $\operatorname{spt}_{d_{\infty}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})=X$ and $\operatorname{spt}_{d_{Z}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})=Z_{0}$.
Finally, we claim the masses converge as in (4.86). By Ambrosio-Kirchheim’s semicontinuity of mass combined with the fact that $\zeta_{\infty}$ is distance preserving we have
(4.106) ${\mathbf{M}}_{d_{\infty}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})={\mathbf{M}}_{d_{Z}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{T}_{\infty}})\leq\liminf_{k\to\infty}{\mathbf{M}}_{d_{Z}}(T_{j_{k}})=\liminf_{k\to\infty}{\mathbf{M}}_{d_{j_{k}}}(T).$
Applying Lemma 4.7, and then the fact that “${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}}=T$ as integral currents on $(X,d_{j_{k}})$”, we have
(4.107) ${\mathbf{M}}_{d_{\infty}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})\geq{\mathbf{M}}_{d_{j_{k}}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{\infty}})={\mathbf{M}}_{d_{j_{k}}}(T).$
Combining (4.106) and (4.107), we have the claimed mass convergence for the subsequence in (4.86). ∎
4.10. Proving SWIF Convergence
In this section we complete the proof of Theorem 4.16. Recall Definition 4.15 of the SWIF distance. Our goal is to provide a constructive proof that
(4.108) $d_{SWIF}((X,d_{j},T),(M_{\infty},d_{\infty},T_{\infty}))\to 0\textrm{ as }j\to\infty$
where $(M_{\infty},d_{\infty},T_{\infty})$ is the integral current space defined using a subsequence in Proposition 4.17.
To complete the proof of Theorem 4.16 we will explicitly estimate the SWIF distances. First we prove a more general proposition that provides a more general estimate on the SWIF distance between a pair of integral current spaces:
Proposition 4.18.
Given two metric spaces, $(X,d_{a})$ and $(X,d_{b})$, and an $\epsilon>0$ such that
(4.109) $d_{b}(x,y)-\epsilon\leq d_{a}(x,y)\leq d_{b}(x,y)\quad\forall x,y\in X.$
If $T_{a}$ is an integral current on $(X,d_{a})$ and $T_{b}$ is an integral current on $(X,d_{b})$ both of dimension, $m$, and “$T_{a}=T_{b}$ as integral currents on $(X,d_{a})$”, then
(4.110) $d_{SWIF}((M_{a},d_{a},T_{a}),(M_{b},d_{b},T_{b})\leq h2^{\frac{m+1}{2}}({\mathbf{M}}_{d_{b}}(T_{b})+{\mathbf{M}}_{d_{b}}(\partial T_{b}))$
where $M_{a}=\operatorname{set}(T_{a})$, $M_{b}=\operatorname{set}(T_{b})$, and $h=\epsilon/2$.